# Differential Test Performance and Peer Effects

### v1

@@ Appendix C Proofs
observing that $\rho_{0}-\tilde{\rho}=o_{p}\left(1\right)$ by Theorem 4.1(ii), that the first and second supremum converges to zero a.s. by Assumption 9(i) and that $\lim_{n}E\left[n^{-1}\operatorname{tr}\left(C\left(\rho\right)\Omega_{0}\right)\right]$ is uniformly bounded by Assumption 9(ii). Arguments analogous to these show that the remaining terms of $n^{-1}\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1(\tau_{c}=j)$ converge to zero in probability. The details are omitted. This establishes that
@@
where $\left\|\bar{\theta}-\theta_{0}\right\|\leq\left\|\hat{\theta}-\theta_{0}\right\|$ and with some abuse of notation it is understood that the between value $\bar{\theta}$ differs among rows of $m_{n}\left(.\right).$ Using the explicit derivatives given in Appendix D it follows from Assumption 9(i) and (ii) that for any sequence $\gamma_{n}$ with $\gamma_{n}\rightarrow_{p}\gamma_{*}$ we have $n^{-1/2}\partial m_{n}\left(\hat{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ and $n^{-1/2}\partial m_{n}\left(\bar{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ where $G(\gamma_{\ast})$ is fixed matrix with full column rank and $\hat{\Xi}_{n}\rightarrow_{p}\Xi$ with $\Xi$ a full-rank square matrix. This implies that
@@
The moment bounds of A-KP 1(i) hold by Assumption 2 and Lemma 2.1 which implies that $u_{it}$ and $\mu_{i}^{\ast}$ have bounded moments, and Assumption 7(i) which bounds the moments of $x_{it}$ and $z_{it}.$ A-KP 1(ii), Eq. (15), holds by Lemma B.3, A-KP 1(ii), Eq. (16) holds by Assumption 3 and A-KP 1(ii), Eq. (17) holds by Lemma B.3, and the parametric restrictions on the conditional cross-sectional variances hold by Assumption 3.
Now turning to A-KP 2, the $\sigma$-field that corresponds to $\mathcal{B}_{n,t}$ is $\mathcal{A}_{n,\left(t-1\right)n}$ defined in (43). Since $\mathcal{Z}_{n}^{*}\subset\mathcal{A}_{n,\left(t-1\right)n}$, the fact that $h_{i}$ and $a_{ij}$ are assumed measurable w.r.t to $\mathcal{Z}_{n}\subset\mathcal{Z}_{n}^{*}$ and elements of $h_{i}$ and the row sums of the absolute elements (and powers thereof) of $A$ are uniformly bounded by Assumptions 5(i)(iii) and 7, A-KP 2(i) and (ii) hold. We also normalize $f_{2}=1$ and restrict the parameter space of $f_{1}$ to a compact interval in Assumption 4 such that A-KP 2(iii) holds.
@@
Appendix D Gradient Vector
In this section, we derive the gradient vector for the two-step-GMM estimator, which is necessary for the limiting variance covariance matrix in Theorem 4.2. Recall that $\theta=\left(\rho,f_{1},\delta^{\prime}\right)^{\prime}$,
@@ Appendix D Gradient Vector
It thus is obvious that $u^{+}(\theta,\gamma)$ and each element of $\frac{u^{+}(\theta,\gamma)}{\partial\theta}$ can be written as linear combinations of $S(\theta,\gamma)\Upsilon$, where $S(\theta,\gamma)$ is an element or the product of elements of the matrix set in (42) and $\Upsilon\in\left\{H,\mu^{*},u_{1},u_{2}\right\}$. Applying Corollary B.1 and Assumption 9, the gradient vector converges uniformly to its expected value.
Appendix E GMM with Randomly Missing Variables
@@
Suppose class $c$ has $n_{c}$ students, and $n_{c}^{obs}\leqslant n_{c}$ of them are observed, which we define as having non-missing values for all variables. Without loss of generality, we order the vectors for the class such that observed students are at the top and unobserved are at the bottom. That is, $y_{c1}=(y_{c1}^{obs\prime},y_{c1}^{unobs\prime})^{\prime}$, where $y_{c1}$ is the $n_{c}\times 1$ vector of test score 1 for all students in class $c$, $y_{c1}^{obs}$ is the $n_{c}^{obs}\times 1$ vector of test score 1 for all observed students, and $y_{c1}^{unobs}$ is the $n_{c}^{unobs}\times 1$ vector of test score 1 for all unobserved students. Note that $n_{c}^{unobs}=n_{c}-n_{c}^{obs}$. Similarly, $y_{c2}=(y_{c2}^{obs\prime},y_{c2}^{unobs\prime})^{\prime}$, $x_{c}=(x_{c}^{obs\prime},x_{c}^{unobs\prime})$, etc.
@@
$y_{1c}-f_{1}y_{2c}=\text{$\underline{X}$}_{c}\delta+(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}).$ (82)
[…]
@@
where $\bar{x}_{c}^{p}=\frac{1}{n_{c}}\sum_{r=1}^{n_{c}}x_{rc}^{p}$ is the average characteristics of all students. Let $\bar{x}_{c}^{p,obs}$ be the averages of all observed $x$. That is, as long as a variable in $x^{p}$ is observed for a student, it will be used to calculate the average even if the student has other variables missing and counted as unobserved. For example, some classes may have variable “$girl$” observed for all students, while race and test information may be missing for some. In this way we increase the efficiency of our approximation as having missing value for a variable is much less common than having missing values for all the variables. Then $\bar{girl}_{c}^{p,obs}=\bar{girl}_{c}^{p}$ as we use all observed values of "$girl"$. Assuming that observations are missing at random it follows that $E\left[\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs}\right]=0$ and the difference between the two is very small as personal characteristics don’t have that many missing values. See Table 9 for details. We can therefore approximate $D_{c}M_{c}x_{c}^{p}$ with $\tilde{x}_{c}^{m,obs}=\frac{1}{n_{c}-1}(\iota_{n_{c}^{obs}}n_{c}\bar{x}_{c}^{p,obs}-x_{c}^{p,obs})$. In all,
$y_{1c}^{obs}-f_{1}y_{2c}^{obs}$ $=x_{c}^{c,obs}\delta_{c}+x_{c}^{p,obs}\delta_{p}+\tilde{x}_{c}^{m,obs}\delta_{m}+v_{c},$
[…]
@@
$\Omega^{obs}\equiv E\left[\epsilon_{c}\epsilon_{c}^{\prime}\right]=\sigma_{c}^{2}\left((\frac{n_{c}-1-\rho}{n_{c}-1})^{2}I_{n^{obs}}+\frac{1}{n_{c}}((1+\rho)^{2}-(\frac{n_{c}-1-\rho}{n_{c}-1})^{2})\iota_{n^{obs}}\iota_{n^{obs}}^{\prime}\right).$ (85)
Note that $E\left[v_{c}|\text{$\underline{x}$}_{c}^{obs}\right]$ and $\textrm{var}(v_{c}|\text{$\underline{x}$}_{c}^{obs})$ requires additional assumptions on the distribution of $\xi_{c}$, in particular $E\left[\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs}|\bar{x}_{c}^{p,obs}\right]=0$. In practice, we ignore the term $\xi_{c}$. As long as the number of missing $x^{p}$ is small, as Table 9 shows, the bias is small and negligible.
@@
$u_{c}^{+obs}=(\Omega^{obs})^{-1/2}\left(y_{1c}^{obs}-f_{1}y_{2c}^{obs}-\text{$\underline{x}$}_{c}^{obs}\delta\right)$
and then continue with our GMM estimation strategy.
Graham(2008) discusses missing values in the online appendix. For individual characteristics, “Twenty three kindergarten student records were missing information on free and reduced price school lunch eligibility, in these cases the missing values were replaced with either eligibility status for the same student in the closest of first, second or third grade (17 cases) or the median value among kindergarten students in their school (6 cases). In three cases missing student race values were replaced with school median values.” For test scores, he just adjust for the group size like what we do above “The analysis below assumes the pattern of missing test score data is indeed random and hence ignorable. Observed test scores are regarded as a random sample from different classrooms of known size. ”

### v2

@@ Appendix D Asymptotic Normality of Parameter Estimators
observing that $\rho_{0}-\tilde{\rho}=o_{p}\left(1\right)$ by Theorem C.1(ii), that the first and second supremum converges to zero a.s. by Assumption 11(i) and that $\lim_{n}E\left[n^{-1}\operatorname{tr}\left(C\left(\rho\right)\Omega_{0}\right)\right]$ is uniformly bounded by Assumption 11(ii). Arguments analogous to these show that the remaining terms of $n^{-1}\sum_{c=1}^{C}\tilde{\epsilon}_{c}^{\prime}\tilde{\epsilon}_{c}1(\tau_{c}=j)$ converge to zero in probability. The details are omitted. This establishes that
@@
where $\left\|\bar{\theta}-\theta_{0}\right\|\leq\left\|\hat{\theta}-\theta_{0}\right\|$ and with some abuse of notation it is understood that the between value $\bar{\theta}$ differs among rows of $m_{n}\left(.\right).$ Using the explicit derivatives given in Appendix F it follows from Assumption 11(i) and (ii) that for any sequence $\gamma_{n}$ with $\gamma_{n}\rightarrow_{p}\gamma_{*}$ we have $n^{-1/2}\partial m_{n}\left(\hat{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ and $n^{-1/2}\partial m_{n}\left(\bar{\theta},\gamma_{n}\right)/\partial\theta\rightarrow_{p}G(\gamma_{\ast})$ where $G(\gamma_{\ast})$ is fixed matrix with full column rank and $\hat{\Xi}_{n}\rightarrow_{p}\Xi$ with $\Xi$ a full-rank square matrix. This implies that
@@
The moment bounds of A-KP 1(i) hold by Assumption 2 and Theorem 2.1 which implies that $u_{it}$ and $\mu_{i}^{\ast}$ have bounded moments, and Assumption 8(i) which bounds the moments of $x_{it}$ and $z_{it}.$ A-KP 1(ii), Eq. (15), holds by Lemma B.1, A-KP 1(ii), Eq. (16) holds by Assumption 9 and A-KP 1(ii), Eq. (17) holds by Lemma B.1, and the parametric restrictions on the conditional cross-sectional variances hold by Assumption 9.
Now turning to A-KP 2, the $\sigma$-field that corresponds to $\mathcal{B}_{n,t}$ is $\mathcal{A}_{n,\left(t-1\right)n}$ defined in (45). Since $\mathcal{Z}_{n}^{*}\subset\mathcal{A}_{n,\left(t-1\right)n}$, the fact that $h_{i}$ and $a_{ij}$ are assumed measurable w.r.t to $\mathcal{Z}_{n}\subset\mathcal{Z}_{n}^{*}$ and elements of $h_{i}$ and the row sums of the absolute elements (and powers thereof) of $A$ are uniformly bounded by Assumptions 6(i)(iii) and 8, A-KP 2(i) and (ii) hold. We also normalize $f_{2}=1$ and restrict the parameter space of $f_{1}$ to a compact interval in Assumption 5 such that A-KP 2(iii) holds.
@@
Appendix E Auxiliary Lemmas
To prove Theorems C.1 and D.1, we establish two preliminary lemmas. We also provide proofs for Lemma 5.1 and Corollary 5.1 in this section. Note that many matrices in this paper, e.g., $M$, $\Omega(\gamma)$, $\Sigma_{t}$ defined in Assumption 9 can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$where $I_{c}^{\ast}=I_{c}-\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}/n_{c}$, $J_{c}^{\ast}=\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}/n_{c}$ are the residual projection matrix and the projection matrix onto $\mathbf{1}_{c}$ respectively. To see this note that
Lemma E.1.
(vi) $\operatorname{tr}(pI_{c}^{\ast}+qJ_{c}^{\ast})=p(n_{c}-1)+q$.
Corollary E.1.
Suppose Assumptions 5, 6, 7, 8, 9, and 12 hold. Then $I+\rho M$, $\Omega(\gamma)$, and $\Sigma_{t}$ are non-singular. If $S(\theta,\gamma)$ is an element or product matrix of elements from
Proof.
By Assumptions 5, 7 and 9, $\left(n_{c}-1-\rho\right)/\left(n_{c}-1\right)$, $\left(1+\rho\right)$, $\gamma_{j}^{2}$, $\rho_{0,\tau_{c}}^{2}$, and $\sigma_{0,t}^{2}$ in (73), (74), (75) are positive and uniformly bounded away from 0. Hence by Lemma E.1(v), $I+\rho M$, $\Omega(\gamma)$ and $\Sigma_{t}$ are non-singular. An inspection of Equations (72)-(75) shows that all matrices in the set can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$ with $p_{c}$ and $q_{c}$ continuously differentiable in $\theta$ and $\gamma$ and uniformly bounded. Also note that by Assumption 6, the row and column sums of $A$ are uniformly bounded in absolute value. Utilizing remark A.1 in Kelejian and Prucha (2004) the corollary thus follows from Lemma E.1. ∎
Lemma E.2.
Suppose Assumptions 5, 6, 7 and 9 hold. Consider the block diagonal matrix $\mathcal{M}\left(\rho\right)=\operatorname{diag}{}_{c=1}^{C}\left(\mathcal{M}_{c}\left(\rho\right)\right)$ with $\mathcal{M}_{c}\left(\rho\right)=\left(I_{c}+\rho M_{c}\right)^{-1}\left(I_{c}+\rho_{0}M_{c}\right)$. Then for all $\gamma\in\Gamma$,
and
Proof.
Noting that $\operatorname{tr}\left(A_{c}\right)=0$ and hence $\operatorname{tr}\left(A_{c}I_{c}^{\ast}\right)=\operatorname{tr}\left(A_{c}(I-J_{c}^{\ast})\right)=-\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$, and in light of Lemma E.1
We have
where
Hence
Next observe that $\operatorname{tr}\left(A_{c}J_{c}^{*}\right)=\mathbf{1}_{c}^{{}^{\prime}}A_{c}\mathbf{1}_{c}/n_{c}\geqslant\text{\text@underline{K}}_{a}>0$ and that for any $n_{c}\geqslant 2$ and $-1<\rho<1$, both $n_{c}-1-\rho$ and $1+\rho$ are positive, and therefore $C(\rho,n_{c})$ is positive for all $\rho\in\left(-1,1\right)$. Therefore, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]<0$ if $\rho>\rho_{0}$, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]>0$ if $\rho<\rho_{0}$ and $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]$ has the same sign for all $c$. In all, for any value of $\gamma$ the only solution for
is $\rho=\rho_{0}$. Furthermore, by Assumptions 5 and 7 there exists some constants $c_{D}$ and $c_{\gamma}$ such that for any $\rho_{0}$ in the interior of the parameter space, $C(\rho,n_{c})>c_{D}>0$ and $\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}>c_{\gamma}>0$. This implies that,
$\left|\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\sum_{c=1}^{C}\text{\text@underline{K}}_{a}c_{D}c_{\gamma},$
$\left|\frac{1}{n}\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\text{\text@underline{K}}_{a}c_{D}c_{\gamma}\frac{C}{n}.$
Proofs of Additional Identification Results.
Proof of Lemma 5.1 and Corollary 5.1.
Lemma 5.1 is a special case of Theorem 5.1 without covariates. Note that Assumptions 3, 4, 5, 6, 7 and 9 hold. Without covariates Equation (42) reduces to $K_{n}\left(\rho,\gamma\right)=z^{\prime}V\left(\rho,\gamma\right)\mu^{*}$ where $V\left(\rho,\gamma\right)=\Omega\left(\gamma\right)^{-1/2}\left(I+\rho M\right)^{-1}.$
Recall that $z=\left[z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right]^{\prime}$ with $z_{c}=\grave{z}_{c}\mathbf{1}_{c}$. Utilizing Lemma E.1,
leads to
where the inequality follows from the conditions imposed in the lemma, in particular (24). With $\rho\in(-1,1)$, $E\left[K_{n}\left(\rho,\gamma\right)|\mathcal{Z}_{n}\right]\neq 0$ such that (38) only has one solution, $f_{1}=f_{1,0}$. Identification of $\rho_{0}$ from the quadratic moment is not affected by the absence of covariates. Thus Lemma 5.1 follows from Theorem 5.1.
Appendix F Gradient Vector
In this section, we derive the gradient vector for the two-step-GMM estimator, which is a component of the limiting variance covariance matrix in Theorem D.1. Recall that $\theta=\left(\rho,f_{1},\delta^{\prime}\right)^{\prime}$,
@@ Appendix F Gradient Vector
It thus is obvious that $u^{+}(\theta,\gamma)$ and each element of $\frac{u^{+}(\theta,\gamma)}{\partial\theta}$ can be written as linear combinations of $S(\theta,\gamma)\Upsilon$, where $S(\theta,\gamma)$ is an element or the product of elements of the matrix set in (76) and $\Upsilon\in\left\{H,\mu^{*},u_{1},u_{2}\right\}$. Applying Corollary E.1 and Assumption 11, the gradient vector converges uniformly to its expected value.
Appendix G GMM with Randomly Missing Variables
@@
Suppose class $c$ has $n_{c}$ students, and $n_{c}^{obs}\leqslant n_{c}$ of them are observed. which We define the case where $n_{c}^{obs}=n_{c}$ as having non-missing values for all variables. Without loss of generality, we order the vectors for the class such that observed students are at the top and unobserved are at the bottom. That is, $y_{c1}=(y_{c1}^{obs\prime},y_{c1}^{unobs\prime})^{\prime}$, where $y_{c1}$ is the $n_{c}\times 1$ vector of test score 1 for all students in class $c$, $y_{c1}^{obs}$ is the $n_{c}^{obs}\times 1$ vector of test score 1 for all observed students, and $y_{c1}^{unobs}$ is the $n_{c}^{unobs}\times 1$ vector of test score 1 for all unobserved students. Note that $n_{c}^{unobs}=n_{c}-n_{c}^{obs}$. Similarly, $y_{c2}=(y_{c2}^{obs\prime},y_{c2}^{unobs\prime})^{\prime}$, $x_{c}=(x_{c}^{obs\prime},x_{c}^{unobs\prime})^{\prime}$, etc.
@@
$y_{1c}-f_{1}y_{2c}=\text{\text@underline{X}}_{c}\delta+(I_{c}+\rho M_{c})(u_{1c}-f_{1}u_{2c}).$ (77)
where $\underline{X}_{c}=\left(v_{c}^{c},w_{1,c}^{c},w_{2,c}^{c},v_{c}^{p},w_{1,c}^{p},w_{2,c}^{p},M_{c}v_{c}^{p},M_{c}w_{1,c}^{p},M_{c}w_{2,c}^{p}\right)$. Let $x_{c}^{c}=[v_{c}^{c},w_{1,c}^{c},w_{2,c}^{c}]$ be class level variables, $x_{c}^{p}=[v_{c}^{p},w_{1,c}^{p},w_{2,c}^{p}]$ be student personal level characteristics. The model for class $c$ then can be written more explicitly as
@@
where $\bar{x}_{c}^{p}=\frac{1}{n_{c}}\sum_{r=1}^{n_{c}}x_{rc}^{p}$ is the average characteristics of all students. Let $\bar{x}_{c}^{p,obs}$ be the averages of all observed $x$. That is, as long as a variable in $x^{p}$ is observed for a student, it will be used to calculate the average even if the student has other variables missing and counted as unobserved. For example, some classes may have variable “$girl$” observed for all students, while race and test information may be missing for some. In this way we increase the efficiency of our approximation as observing a missing value for a particular variable is much less common than having missing values for all of the variables related to a particular student. Then $\bar{girl}_{c}^{p,obs}=\bar{girl}_{c}^{p}$ as we use all observed values of "$girl"$. Assuming that observations are missing at random it follows that $E\left[\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs}\right]=0$ and the difference between the two is very small as personal characteristics don’t have that many missing values. See Table 4 for details. We can therefore approximate $D_{c}M_{c}x_{c}^{p}$ with $\tilde{x}_{c}^{m,obs}=\frac{1}{n_{c}-1}(\iota_{n_{c}^{obs}}n_{c}\bar{x}_{c}^{p,obs}-x_{c}^{p,obs})$. In all,
$y_{1c}^{obs}-f_{1}y_{2c}^{obs}$ $=x_{c}^{c,obs}\delta_{c}+x_{c}^{p,obs}\delta_{p}+\tilde{x}_{c}^{m,obs}\delta_{m}+v_{c}=\text{\text@underline{x}}_{c}^{obs}\delta+v_{c},$
@@
$\Omega_{c}^{obs}\equiv E\left[\epsilon_{c}\epsilon_{c}^{\prime}\right]=\sigma_{c}^{2}\left((\frac{n_{c}-1-\rho}{n_{c}-1})^{2}I_{n^{obs}}+\frac{1}{n_{c}}((1+\rho)^{2}-(\frac{n_{c}-1-\rho}{n_{c}-1})^{2})\iota_{n^{obs}}\iota_{n^{obs}}^{\prime}\right).$ (80)
Note that $E\left[v_{c}|\text{\text@underline{x}}_{c}^{obs}\right]$ and $\textrm{var}(v_{c}|\text{\text@underline{x}}_{c}^{obs})$ requires additional assumptions on the distribution of $\xi_{c}$, in particular $E\left[\bar{x}_{c}^{p}-\bar{x}_{c}^{p,obs}|\bar{x}_{c}^{p,obs}\right]=0$. In practice, we ignore the term $\xi_{c}$. As long as the number of missing $x^{p}$ is small, as Table 4 shows, the bias is small and negligible.
@@
$u_{c}^{+obs}=(\Omega_{c}^{obs})^{-1/2}\left(y_{1c}^{obs}-f_{1}y_{2c}^{obs}-\text{\text@underline{x}}_{c}^{obs}\delta\right)$
and continue with our GMM estimation strategy.
Appendix H Additional Empirical Results
H.1 Additional Specification Tests
Our analysis thus far has documented that the word and reading scores are highly correlated both contemporaneously as well as over the course of one school year. This evidence is consistent with restrictions imposed by Assumption 2 and the hypothesis that closely related scores depend on the same unobserved components. A caveat is that closely related scores may remove too much common variation when quasi-differences of these scores are used and thus render statistical evidence about peer effects imprecise. Panel A of Table 8 reports pseudo $R^{2}$ measures of model estimates for word and reading scores and model specifications (1) through (4) with column (1) only controlling for school fixed effects, and column (4) being the full specification with school and class type effects, student and teacher characteristics as well as peer characteristics. Results for other pairs of scores are presented in panels B-F as a comparison.
The observed pseudo $R^{2}$ measures are all below 0.2 for the word-read pair. As expected the pseudo $R^{2}$ generally increases with larger models, or in other words by moving from Column (1) to (4). Compared to other pairs of scores, the word-read score pair often has the lowest $R^{2}$ which is an indication that the differential score in these specifications already controls for most of the variation explained by the covariates. We take this as further evidence that word-read best fits the model assumptions we impose.
Table 9 presents standard deviations of $Q_{x}y$ for reading and word scores and $Q_{x}(y_{1}-f_{1}y_{2})$ for $y_{1}$ being word and $y_{2}$ being reading scores for Grade K to 3. The left and right panels are for small and regular classes respectively. Control variables for specifications (1)-(4) are the same as those in Table 8, as discussed above. We see that in Grade K to 3, the ratio of standard deviations of $Q_{x}(y_{1}-f_{1}y_{2})$ to that of $Q_{x}y$ is 1/5 to 1/4.This shows that quasi-differencing removes a significant part of the variation in the test scores, consistent with the term $\mu^{*}$ playing a major role in determining test performance. At the same time, the results in Table 9 indicate that the inclusion of controls does not eliminate much of the remaining variation not captured by fixed effects. The conclusion is that after differencing there is between 20 to 30 percent of test score variation remaining for the identification of $\rho$.
H.2 Effects of Individual Characteristics
We also provide the coefficients of peer averages of individual characteristics (age, black, free lunch, and girl) in Table 10. These four variables all pertain to test-invariant individual characteristics. The estimated coefficients can be expressed as $\rho(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})$, where $\rho\beta_{v1}^{p}$ and $\rho\beta_{v2}^{p}$ represent the exogenous peer effect of the variable on tests $t=1$ and $t=2$ respectively. Thus, these coefficients capture the differential exogenous peer effects across the two tests. We also scale the coefficients for dummies $girl$, $black$ and $free\,lunch$ by 1/100 so that the coefficients are the impact on the outcome of 1 percentage point changes in share of peers who are girl/black/free lunch receivers. Except for kindergarten where the presence of girls among peers positively impacts reading more than word study, none of the four variables show significant differential exogenous peer effects. Overall, the results indicate small and mostly insignificant peer effects on differential scores. This is consistent with a scenario where $y_{2}$ is essentially sufficient to control for both observed and unobserved test invariant determinants of test performance. In fact, when $f_{1}=1$ and $\beta_{v1}^{p}=\beta_{v2}^{p}$ these effects would be zero in our model and the overall evidence points to the conclusion that this is approximately the case in the sample as well.